# bibliography-2.bib

### v5

@@ 2.2. Real embedded Grassmannians have projectively ${\mathbb{Q}}$-algebraic homology
In [GSa, Definition 2.23], Ghiloni and the author introduced a variant ‘over ${\mathbb{Q}}$’ of the classical algebraic representatives of homology classes of a topological space. Here we briefly recall the notion of ‘projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-algebraic homology’. Let $W\subset{\mathbb{R}}^{k}$ be a set endowed with the induced Euclidean topology. Given $p\in{\mathbb{N}}$ and $\alpha\in H_{p}(W,{\mathbb{Z}}/2{\mathbb{Z}})$, we say that $\alpha$ is projectively ${\mathbb{Q}}$-algebraic if there exist a $p$-dimensional projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic set $Z\subset{\mathbb{R}}^{h}$ and a ${\mathbb{Q}}$-regular map $g:Z\rightarrow W$ such that $g_{*}([Z])=\alpha$, where $[Z]$ is the fundamental class of $Z$ in $H_{p}(Z,{\mathbb{Z}}/2{\mathbb{Z}})$.
@@
By classical arguments about cellular and simplicial homology, $\big{\{}[\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big{\}}$ constitutes a system of generators of $H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})$ for every $k=0,\dots,d$. We are only left to prove that $\big{\{}[\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big{\}}$ is linearly independent over ${\mathbb{Z}}/2{\mathbb{Z}}$. Since $\overline{\Omega}$ is algebraic for every open cell $\Omega\in\mathcal{S}^{(k)}$, for every $k\in\{0,\dots,d\}$, the fundamental class $[\overline{\Omega}]$ of $\overline{\Omega}$ is a well defined homology class in $H_{k}(W,{\mathbb{Z}}/2{\mathbb{Z}})$, see [BCR98, § 11.3]. Suppose $\Omega\in\mathcal{S}^{(k)}$, then for every $\Omega^{\prime}\in\mathcal{S}^{(k+1)}$ we have $\partial_{k+1}^{\mathcal{S}}(\overline{\Omega}^{\prime})=0$, since $\overline{\Omega}^{\prime}$ is algebraic as well. Hence, we get that $[\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})$ is non-null and linearly independent with respect to $\big{\{}[\overline{\Omega}^{\prime}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega^{\prime}\in\mathcal{S}^{(k)}\,\text{and}\,\Omega^{\prime}\neq\Omega\big{\}}$ for every choice of $\Omega\in\mathcal{S}^{(k)}$ and $k\in\{0,\dots,d\}$. This proves that $\big{\{}[\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big{\}}$ is a basis of $H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})$, then $\big{\{}[\overline{\Omega}]\in H_{k}(W,{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big{\}}$ it is also a basis of $H_{k}(W,{\mathbb{Z}}/2{\mathbb{Z}})$, as desired. ∎
@@ 3.1. ${\mathbb{Q}}$-Algebraic bordism classes and unoriented relative bordisms
In [GSa], Ghiloni and the author introduced a variant ‘over ${\mathbb{Q}}$’ of the classical algebraic unoriented bordism and investigated its relation with projectively ${\mathbb{Q}}$-algebraic homology. Here we briefly recall those notions and useful results. Let $W\subset{\mathbb{R}}^{k}$ be a real algebraic set. Given a compact $\mathscr{C}^{\infty}$ manifold $P$ and a $\mathscr{C}^{\infty}$ map $f:P\to W$, we say that the unoriented bordism class of $f$ is projectively ${\mathbb{Q}}$-algebraic if there exist a compact $\mathscr{C}^{\infty}$ manifold $T$ with boundary, a projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic set $Y\subset{\mathbb{R}}^{h}$, a $\mathscr{C}^{\infty}$ diffeomorphism $\psi:P\sqcup Y\to\partial T$ and a $\mathscr{C}^{\infty}$ map $F:T\rightarrow W$ such that $F\circ\jmath\circ(\psi|_{P})=f$ and $F\circ\jmath\circ(\psi|_{Y})$ is a ${\mathbb{Q}}$-regular map, where $\jmath:\partial T\hookrightarrow T$ is the inclusion map.
@@
Here we recall the fundamental result of [GSa] about the equivalence between Definitions 2.9 & 3.1. That is:
Theorem 3.2 ([GSa, Theorem 2.25]).
@@
Let us specify ‘over ${\mathbb{Q}}$’ the construction of the algebraic unoriented relative cobordism by Akbulut and King in [AK81b, Lemma 4.1].
@@
$S^{\alpha}$ :=$:=\{(x,y_{1},t_{1},\dots,y_{|\alpha|},t_{|\alpha|})\in T_{\alpha}\times({\mathbb{R}}^{n}\times{\mathbb{R}})^{|\alpha|}\,|\,(\mu^{\alpha}(x),y_{1},t_{1},\dots,y_{|\alpha|},t_{|\alpha|})\in{\mathbb{E}}_{\alpha}^{*}\}$
$S_{i}^{\alpha}$ :=$:=\{(x,y_{1},t_{1},\dots,y_{|\alpha|},t_{|\alpha|})\in S^{\alpha}\,|\,y_{i}=0,\,t_{i}=0\},$
@@
$N_{i}$ :=$:=\begin{cases}N\cap S_{i}\quad&\text{if $i\in\alpha$,}\\ M_{i}\times\{1\}&\text{otherwise.}\end{cases}$
@@
Define $T:=S\cup T^{\prime}$ and $T_{i}:=S_{i}\cup T^{\prime}_{i}$, after smoothing corners. Let $k:=\max\{k_{\alpha},k^{\prime}\}$ and consider $\iota_{\alpha}:{\mathbb{R}}^{k_{\alpha}}\to{\mathbb{R}}^{k}$ and $\iota^{\prime}:{\mathbb{R}}^{k^{\prime}}\to{\mathbb{R}}^{k}$ be the inclusion mappings. Then, after a translation of a rational factor $v\in{\mathbb{Q}}^{k}$ if necessary, we may assume that $(\iota^{\prime}(Y^{\prime})+v)\cap\iota_{\alpha}(Y^{\alpha})=\varnothing$, thus $Y:=\iota_{\alpha}(Y^{\alpha})\sqcup(\iota^{\prime}(Y^{\prime})+v)\subset{\mathbb{R}}^{k}$ is a projectively ${\mathbb{Q}}$-closed ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic set by [GSa, Lemmas 2.11(b) & 2.15(a)(c)]. Let $\psi:M\sqcup Y\to\partial T$ defined as follows $\psi|_{M}:=\psi_{\alpha}|_{M}$, $\psi|_{\iota_{\alpha}(Y_{\alpha})}(x):=\psi_{\alpha}(\iota_{\alpha}^{-1}(x))$ and $\psi|_{\iota^{\prime}(Y^{\prime})+v}(x):=\psi^{\prime}((\iota^{\prime})^{-1}(x-v))$. ∎
@@
Let us construct the desired compact $\mathscr{C}^{\infty}$ submanifold $S$ of ${\mathbb{R}}^{s+1}={\mathbb{R}}^{s}\times{\mathbb{R}}$, following the strategy used by Tognoli in [Tog73, § b), pp. 176-177]. By the collaring theorem (see [Hir94, Theorem 6.1, p. 113]), there exist an open neighborhood $U$ of $\partial T$ in $T$ and a $\mathscr{C}^{\infty}$ diffeomorphism $\phi^{\prime}:U\to\partial T\times[0,1)$ such that $\phi^{\prime}(t)=(t,0)$ for all $t\in\partial T$ and $\phi^{\prime}|_{T_{i}\cap U}:T_{i}\cap U\to\partial T_{i}\times[0,1)$ is a diffeomorphism as well, for every $i=1,\dots,\ell$. Let $\phi:U\to(M\sqcup Y)\times[0,1)$ be the $\mathscr{C}^{\infty}$ diffeomorphism $\phi:=(\psi^{-1}\times\mathrm{id}_{[0,1)})\circ\phi^{\prime}$. Note that $\phi(t)=(\psi^{-1}(t),0)$ for all $t\in\partial T$. Set $A:=T\setminus\partial T$, $B:=\phi^{-1}((M\sqcup Y)\times(0,\frac{1}{2}])\subset A$, $N:={\mathbb{R}}^{s}\times(0,+\infty)$ and define the map $\theta:B\to N$ by $\theta(x,x_{s+1}):=\phi(x,x_{s+1})$. Since we can safely assume $s+1\geq 2(d+1)+1$, Tietze’s theorem ensures the existence of a continuous extension of $\theta$ from the whole $A$ to $N$, we can apply to $\theta$ the extension theorem [Whi36, Theorem 5$(\mathrm{f})$] of Whitney, obtaining a $\mathscr{C}^{\infty}$ embedding $\Theta:A\to N$ extending $\theta$. Let $R:{\mathbb{R}}^{s+1}={\mathbb{R}}^{s}\times{\mathbb{R}}\to{\mathbb{R}}^{s+1}$ be the reflection $R(x,x_{s+1}):=(x,-x_{s+1})$ and let $S^{\prime}$ and $S^{\prime}_{i}$ be the compact $\mathscr{C}^{\infty}$ submanifolds $\Theta(A)\sqcup((M\sqcup Y)\times\{0\})\sqcup R(\Theta(A))$ and $\Theta(A\cap T_{i})\sqcup((M_{i}\sqcup Y_{i})\times\{0\})\sqcup R(\Theta(A\cap T_{i}))$ of ${\mathbb{R}}^{s+1}$, for every $i=1,\dots,\ell$, respectively. Thanks to the compactness of $T$ and of each $T_{i}$, there exists $\epsilon>0$ such that $S^{\prime}\cap({\mathbb{R}}^{s}\times(-\epsilon,\epsilon))=(M\sqcup Y)\times(-\epsilon,\epsilon)$ and $S^{\prime}_{i}\cap({\mathbb{R}}^{s}\times(-\epsilon,\epsilon))=(M_{i}\sqcup Y_{i})\times(-\epsilon,\epsilon)$. Let $L:{\mathbb{R}}^{s+1}\to{\mathbb{R}}^{s+1}$ be the linear isomorphism $L(x,x_{s+1}):=(x,\epsilon^{-1}x_{s+1})$. The compact $\mathscr{C}^{\infty}$ submanifold $S:=L(S^{\prime})$ with $\mathscr{C}^{\infty}$ submanifolds $S_{i}:=L(S^{\prime}_{i})$, for every $i=1,\dots,\ell$, in general position of ${\mathbb{R}}^{s+1}$ have the desired properties (ii) & (iv). ∎
3.2. A review on ${\mathbb{Q}}$-nice ${\mathbb{Q}}$-algebraic sets
@@
Here we briefly recall the notion of ${\mathbb{Q}}$-nice ${\mathbb{Q}}$-algebraic set and we develop an explicit example of ${\mathbb{Q}}$-nice ${\mathbb{Q}}$-algebraic set useful for applying approximation techniques ‘over ${\mathbb{Q}}$’ in Section 4.
Definition 3.5 ([GSa, Definition 3.1]).
@@
Let $L\subset{\mathbb{R}}^{n}$ be a ${\mathbb{Q}}$-algebraic set. We say that the pair $L$ is a ${\mathbb{Q}}$-nice if for each $a\in L$, there exists an open neighborhood $U_{a}$ of $a$ in ${\mathbb{R}}^{n}$ such that
@@
Evidently, by definition, the disjoint union of finitely many ${\mathbb{Q}}$-nice algebraic subsets of ${\mathbb{R}}^{n}$ is again a ${\mathbb{Q}}$-nice algebraic subset of ${\mathbb{R}}^{n}$. Moreover, by [GSa, Corollary 3.3], every disjoint union of ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic sets is ${\mathbb{Q}}$-nice and, by [GSa, Lemma 3.4], if $L\subset{\mathbb{R}}^{n}$ is a ${\mathbb{Q}}$-nice ${\mathbb{Q}}$-algebraic set then $L\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{k}$ is ${\mathbb{Q}}$-nice for every $k\in{\mathbb{N}}$. Next Lemma 3.6 will be very useful in the setting of Theorem 4.2.
@@
Let $M\subset{\mathbb{R}}^{n}$ be a compact $\mathscr{C}^{\infty}$ manifold of dimension $d$. Let $X\subset M$ be a ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic subset of ${\mathbb{R}}^{n}$ of codimension $c$ and let $Y\subset M$ be a ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic hypersurface of $M$. If the germ $(M,X\cup Y)$ of $M$ at $X\cup Y$ is the germ of a ${\mathbb{Q}}$-nonsingular ${\mathbb{Q}}$-algebraic set, then $X\cup Y$ is ${\mathbb{Q}}$-nice.

### v6

@@ 2.2. Real embedded Grassmannians have projectively $\mathbb{Q}$}-algebraic homology
In [GS25, Definition 2.38], Ghiloni and the author introduced a variant ‘over $\mathbb{Q}$}’ of the classical algebraic representatives of homology classes of a topological space. Here we briefly recall the notion of ‘projectively $\mathbb{Q}$}-closed $\mathbb{Q}$}-algebraic homology’. Let $W\subset{\mathbb{R}}^{k}$ be a set endowed with the induced Euclidean topology. Given $p\in{\mathbb{N}}$ and $\alpha\in H_{p}(W,{\mathbb{Z}}/2{\mathbb{Z}})$, we say that $\alpha$ is projectively $\mathbb{Q}$}-algebraic if there exist a $p$-dimensional projectively $\mathbb{Q}$}-closed $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic set $Z\subset{\mathbb{R}}^{h}$ and a $\mathbb{Q}$}-regular map $g:Z\rightarrow W$ such that $g_{*}([Z])=\alpha$, where $[Z]$ is the fundamental class of $Z$ in $H_{p}(Z,{\mathbb{Z}}/2{\mathbb{Z}})$.
@@
By classical arguments about cellular and simplicial homology, $\big\{[\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big\}$ constitutes a system of generators of $H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})$ for every $k=0,\dots,d$. We are only left to prove that $\big\{[\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big\}$ is linearly independent over ${\mathbb{Z}}/2{\mathbb{Z}}$. Since $\overline{\Omega}$ is algebraic for every open cell $\Omega\in\mathcal{S}^{(k)}$, for every $k\in\{0,\dots,d\}$, the fundamental class $\overline{\Omega}]$ of $\overline{\Omega}$ is a well defined homology class in $H_{k}(W,{\mathbb{Z}}/2{\mathbb{Z}})$, see [BCR98, § 11.3]. Suppose $\Omega\in\mathcal{S}^{(k)}$, then for every $\Omega^{\prime}\in\mathcal{S}^{(k+1)}$ we have $\partial_{k+1}^{\mathcal{S}}(\overline{\Omega}^{\prime})=0$, since $\overline{\Omega}^{\prime}$ is algebraic as well. Hence, we get that $\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})$ is non-null and linearly independent with respect to $\big\{[\overline{\Omega}^{\prime}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega^{\prime}\in\mathcal{S}^{(k)}\,\text{and}\,\Omega^{\prime}\neq\Omega\big\}$ for every choice of $\Omega\in\mathcal{S}^{(k)}$ and $k\in\{0,\dots,d\}$. This proves that $\big\{[\overline{\Omega}]\in H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big\}$ is a basis of $H_{k}(\mathcal{S},{\mathbb{Z}}/2{\mathbb{Z}})$, then $\big\{[\overline{\Omega}]\in H_{k}(W,{\mathbb{Z}}/2{\mathbb{Z}})\,|\,\Omega\in\mathcal{S}^{(k)}\big\}$ it is also a basis of $H_{k}(W,{\mathbb{Z}}/2{\mathbb{Z}})$, as desired. ∎
@@ 3.1. $\mathbb{Q}$}-Algebraic bordism classes and unoriented relative bordisms
In [GS25, Definition 2.39], Ghiloni and the author introduced a variant ‘over $\mathbb{Q}$}’ of the classical algebraic unoriented bordism and investigated its relation with projectively $\mathbb{Q}$}-algebraic homology. Here we briefly recall those notions and useful results. Let $W\subset{\mathbb{R}}^{k}$ be a real algebraic set. Given a compact $\mathscr{C}^{\infty}$ manifold $P$ and a $\mathscr{C}^{\infty}$ map $f:P\to W$, we say that the unoriented bordism class of $f$ is projectively $\mathbb{Q}$}-algebraic if there exist a compact $\mathscr{C}^{\infty}$ manifold $T$ with boundary, a projectively $\mathbb{Q}$}-closed $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic set $Y\subset{\mathbb{R}}^{h}$, a $\mathscr{C}^{\infty}$ diffeomorphism $\psi:P\sqcup Y\to\partial T$ and a $\mathscr{C}^{\infty}$ map $F:T\rightarrow W$ such that $F\circ\jmath\circ(\psi|_{P})=f$ and $F\circ\jmath\circ(\psi|_{Y})$ is a $\mathbb{Q}$}-regular map, where $\jmath:\partial T\hookrightarrow T$ is the inclusion map.
@@
Here we recall the fundamental result of [GS25] about the equivalence between Definitions 2.9 & 3.1. That is:
Lemma 3.2 ([GS25, Lemma 2.41]).
@@
Let us specify ‘over $\mathbb{Q}$}’ the construction of the algebraic unoriented relative cobordism by Akbulut and King in [AK81, Lemma 4.1].
@@
$S^{\alpha}$ $:=\{(x,y_{1},t_{1},\dots,y_{|\alpha|},t_{|\alpha|})\in T_{\alpha}\times({\mathbb{R}}^{n}\times{\mathbb{R}})^{|\alpha|}\,|\,(\mu^{\alpha}(x),y_{1},t_{1},\dots,y_{|\alpha|},t_{|\alpha|})\in{\mathbb{E}}_{\alpha}^{*}\}$
$S_{i}^{\alpha}$ $:=\{(x,y_{1},t_{1},\dots,y_{|\alpha|},t_{|\alpha|})\in S^{\alpha}\,|\,y_{i}=0,\,t_{i}=0\},$
@@
$N_{i}$ $:=\begin{cases}N\cap S_{i}\quad&\text{if $i\in\alpha$,}\\ M_{i}\times\{1\}&\text{otherwise.}\end{cases}$
@@
Define $T:=S\cup T^{\prime}$ and $T_{i}:=S_{i}\cup T^{\prime}_{i}$, after smoothing corners. Let $k:=\max\{k_{\alpha},k^{\prime}\}$ and consider $\iota_{\alpha}:{\mathbb{R}}^{k_{\alpha}}\to{\mathbb{R}}^{k}$ and $\iota^{\prime}:{\mathbb{R}}^{k^{\prime}}\to{\mathbb{R}}^{k}$ be the inclusion mappings. Then, after a translation of a rational factor $v\in{\mathbb{Q}}^{k}$ if necessary, we may assume that $(\iota^{\prime}(Y^{\prime})+v)\cap\iota_{\alpha}(Y^{\alpha})=\varnothing$, thus $Y:=\iota_{\alpha}(Y^{\alpha})\sqcup(\iota^{\prime}(Y^{\prime})+v)\subset{\mathbb{R}}^{k}$ is a projectively $\mathbb{Q}$}-closed $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic set by [GS25, Proposition 2.16(i) & Lemma 2.24]. Let $\psi:M\sqcup Y\to\partial T$ defined as follows $\psi|_{M}:=\psi_{\alpha}|_{M}$, $\psi|_{\iota_{\alpha}(Y_{\alpha})}(x):=\psi_{\alpha}(\iota_{\alpha}^{-1}(x))$ and $\psi|_{\iota^{\prime}(Y^{\prime})+v}(x):=\psi^{\prime}((\iota^{\prime})^{-1}(x-v))$. ∎
@@
Let us construct the desired compact $\mathscr{C}^{\infty}$ submanifold $S$ of ${\mathbb{R}}^{s+1}={\mathbb{R}}^{s}\times{\mathbb{R}}$, following the strategy used by Tognoli in [Tog73, § b), pp. 176-177]. By the collaring theorem (see [Hir94, Theorem 6.1, p. 113]), there exist an open neighborhood $U$ of $\partial T$ in $T$ and a $\mathscr{C}^{\infty}$ diffeomorphism $\phi^{\prime}:U\to\partial T\times[0,1)$ such that $\phi^{\prime}(t)=(t,0)$ for all $t\in\partial T$ and $\phi^{\prime}|_{T_{i}\cap U}:T_{i}\cap U\to\partial T_{i}\times[0,1)$ is a diffeomorphism as well, for every $i=1,\dots,\ell$. Let $\phi:U\to(M\sqcup Y)\times[0,1)$ be the $\mathscr{C}^{\infty}$ diffeomorphism $\phi:=(\psi^{-1}\times\mathrm{id}_{[0,1)})\circ\phi^{\prime}$. Note that $\phi(t)=(\psi^{-1}(t),0)$ for all $t\in\partial T$. Set $A:=T\setminus\partial T$, $B:=\phi^{-1}((M\sqcup Y)\times(0,\frac{1}{2}])\subset A$, $N:={\mathbb{R}}^{s}\times(0,+\infty)$ and define the map $\theta:B\to N$ by $\theta(x,x_{s+1}):=\phi(x,x_{s+1})$. Since we can safely assume $s+1\geq 2(d+1)+1$, Tietze’s theorem ensures the existence of a continuous extension of $\theta$ from the whole $A$ to $N$, we can apply to $\theta$ the extension theorem [Whi36, Theorem 5(f)($\mathrm${f})] of Whitney, obtaining a $\mathscr{C}^{\infty}$ embedding $\Theta:A\to N$ extending $\theta$. Let $R:{\mathbb{R}}^{s+1}={\mathbb{R}}^{s}\times{\mathbb{R}}\to{\mathbb{R}}^{s+1}$ be the reflection $R(x,x_{s+1}):=(x,-x_{s+1})$ and let $S^{\prime}$ and $S^{\prime}_{i}$ be the compact $\mathscr{C}^{\infty}$ submanifolds $\Theta(A)\sqcup((M\sqcup Y)\times\{0\})\sqcup R(\Theta(A))$ and $\Theta(A\cap T_{i})\sqcup((M_{i}\sqcup Y_{i})\times\{0\})\sqcup R(\Theta(A\cap T_{i}))$ of ${\mathbb{R}}^{s+1}$, for every $i=1,\dots,\ell$, respectively. Thanks to the compactness of $T$ and of each $T_{i}$, there exists $\epsilon>0$ such that $S^{\prime}\cap({\mathbb{R}}^{s}\times(-\epsilon,\epsilon))=(M\sqcup Y)\times(-\epsilon,\epsilon)$ and $S^{\prime}_{i}\cap({\mathbb{R}}^{s}\times(-\epsilon,\epsilon))=(M_{i}\sqcup Y_{i})\times(-\epsilon,\epsilon)$. Let $L:{\mathbb{R}}^{s+1}\to{\mathbb{R}}^{s+1}$ be the linear isomorphism $L(x,x_{s+1}):=(x,\epsilon^{-1}x_{s+1})$. The compact $\mathscr{C}^{\infty}$ submanifold $S:=L(S^{\prime})$ with $\mathscr{C}^{\infty}$ submanifolds $S_{i}:=L(S^{\prime}_{i})$, for every $i=1,\dots,\ell$, in general position of ${\mathbb{R}}^{s+1}$ have the desired properties (ii) & (iv). ∎
3.2. A review on $\mathbb{Q}$}-stable $\mathbb{Q}$}-algebraic sets
@@
Here we briefly recall the notion of $\mathbb{Q}$}-stable $\mathbb{Q}$}-algebraic set and we develop an explicit example of $\mathbb{Q}$}-stable $\mathbb{Q}$}-algebraic set useful for applying approximation techniques ‘over $\mathbb{Q}$}’ in Section 4.
Definition 3.5 ([GS25, Definition 3.1]).
@@
Let $L\subset{\mathbb{R}}^{n}$ be a $\mathbb{Q}$}-algebraic set. We say that the pair $L$ is a $\mathbb{Q}$}-stable if for each $a\in L$, there exists an open neighborhood $U_{a}$ of $a$ in ${\mathbb{R}}^{n}$ such that
@@
Evidently, by definition, the disjoint union of finitely many $\mathbb{Q}$}-stable algebraic subsets of ${\mathbb{R}}^{n}$ is again a $\mathbb{Q}$}-stable algebraic subset of ${\mathbb{R}}^{n}$. Moreover, by [GS25, Lemma 3.3(iii)(iv)], every disjoint union of finitely many $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic sets is $\mathbb{Q}$}-stable and, if $L\subset{\mathbb{R}}^{n}$ is a $\mathbb{Q}$}-stable $\mathbb{Q}$}-algebraic set then $L\times\{0\}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}^{k}$ is $\mathbb{Q}$}-stable for every $k\in{\mathbb{N}}$. Next Lemma 3.6 will be very useful in the setting of Theorem 4.2.
@@
Let $M\subset{\mathbb{R}}^{n}$ be a compact $\mathscr{C}^{\infty}$ manifold of dimension $d$. Let $X\subset M$ be a $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic subset of ${\mathbb{R}}^{n}$ of codimension $c$ and let $Y\subset M$ be a $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic hypersurface of $M$. If the germ $(M,X\cup Y)$ of $M$ at $X\cup Y$ is the germ of a $\mathbb{Q}$}-nonsingular $\mathbb{Q}$}-algebraic set, then $X\cup Y$ is $\mathbb{Q}$}-stable.